# EFC++: Elastic Feature Consolidation with Prototype Re-balancing for Cold Start Exemplar-free Incremental Learning

### v2

@@ III-A Exemplar-free Class-Incremental Learning (EFCIL)
III-B Weight Regularization and Feature Distillation
Kirckpatrick et al. [10] showed that using $\ell_{2}$ regularization to constrain weight drift reduces forgetting but does not leave enough plasticity for learning of new tasks. Hence, they proposed Elastic Weight Consolidation (EWC), which relaxes the $\ell_{2}$ constraint with a quadratic constraint based on a diagonal approximation of the Empirical Fisher Information Matrix (E-FIM):
@@ III-B Weight Regularization and Feature Distillation
where $\theta_{t}^{*}$ is the model trained after task $t$. The E-FIM induces a pseudo-metric in parameter space [61] encouraging parameters to stay in a low-error region for previous-task models:
@@
The main limitation of approaches of this type is need for approximations of the E-FIM (e.g. a diagonal assumption). This makes the computation of the E-FIM tractable, but in practice is unrealistic since it does not consider off-diagonal entries that represent influences of interactions between weights on the log-likelihood. To overcome these limitations, more recent EFCIL approaches [21, 22] rely on feature distillation (FD), which scale better in the number of parameters:
$\mathcal{L}_{t}^{\text{FD}}\vcentcolon=\sum_{x\in\mathbf{X}}||f(x;\theta_{t})-f(x;\theta_{t-1})||_{2},$ (5)
@@
The isotropic regularizer using the $\ell_{2}$ distance is too harsh a constraint for learning new tasks [25]. We propose to use a pseudo-metric in feature space, induced by a matrix which we call the Empirical Feature Matrix (EFM), which constrains directions in feature space most important for previous tasks, while allowing more plasticity in other directions when learning new tasks. As shown in Figure 1, increasing plasticity is particularly relevant in Cold Start incremental learning since, in the initial task, a high-quality feature representation cannot be learned and methods must be plastic to adapt their backbone to new tasks.
III-C The Empirical Feature Matrix
Figure 3: The regularizing effects of $E_{t}$ on the Cold Start CIFAR-100 - 10 and 20 step scenarios (see Section VI-A for details on dataset settings). Left: Perturbing features in the principal directions of $E_{1}$ results in significant changes in classifier outputs (in blue), while perturbations in non-principal directions leave the outputs unchanged (in red). Middle: If we continue incremental learning up through task 3 and perturb features from all three tasks in the principal (solid lines) and non-principal (dashed lines) directions of $E_{3}$, we see that $E_{3}$ captures all important directions in feature space up through task 3. Right: At the end of training, we observe the same behavior: the last per-step accuracy (see Eq. 21), representing the average accuracy over all tasks after the last training session, decreases only when perturbed in directions of $E_{10}$ or $E_{20}$ relevant for previous tasks in the 10-step and 20-step scenarios, respectively.
@@ III-C The Empirical Feature Matrix
To gain additional insight into what this matrix measures, we highlight some (information) geometry aspects of $E_{t}$. The Empirical Fisher Information Matrix provides information about the information geometry of parameter space. In particular, it can be obtained as the second derivative of the KL-divergence for small weight perturbations [62, 63]. Similarly, the Empirical Feature Matrix provides information about the information geometry of feature space. Since $E_{f_{t}(x)}$ is positive semi-definite, we can interpret it as a pseudo-metric in feature space and use it to understand which perturbations of features most affect the predicted probability:
@@
where $\mathbf{X}\sim\mathcal{X}_{t}$. In Eq. 10 we employ a damping term $\eta\in\mathbb{R}$ to constrain features to stay in a region where the quadratic second-order approximation of the KL-divergence of the log-likelihood in Eq. LABEL:kl_features remains valid [64]. However, to ensure that our regularizer does not degenerate into feature distillation, we must ensure that $\lambda_{\text{EFM}}\nu_{i}>\eta$, where $\nu_{i}$ are the non-zero eigenvalues of the EFM. Additional analysis of these constraints and the spectrum of the EFM are given in Section VIII-C.
@@
Finally, the rightmost plot in Figure 3 illustrates the per-step accuracy at the end of the training, reflecting the overall continual learning performance (see Section VI-C for a formal definition). As observed earlier, perturbing features along the principal directions of $E_{10}$ (or $E_{20}$) rapidly decreases performance, while perturbations in non-principal directions have no impact on the overall accuracy. This confirms that this behavior remains consistent across all tasks.
IV Elastic Feature Consolidation with Prototype Re-balancing
@@
IV-A Prototype Rehearsal for EFCIL
EFCIL suffers from the fact that the final classifier is never jointly trained on the all seen classes. At each task $t$, the classifier $W_{t-1}$ is extended to include $|\mathcal{C}_{t}|$ new outputs, resulting in a classifier with weights $W_{t}=[W_{t-1},W_{t-1:t}]$. Thus, if classifier $W_{t}$ is only trained on samples from the current classes $\mathcal{C}_{t}$, it becomes incapable of discriminating classes from previous tasks, leading to inter-task confusion. Additionally, the classifier is biased towards the newest classes, resulting in task-recency bias [14, 42].
A common exemplar-free solution is to replay only prototypes of previous classes [20, 21, 22, 23, 26, 55]. At the end of task $t-1$, the prototypes $p^{c}_{t-1}$ are computed as the class means of deep features, and the set of prototypes $\mathcal{P}_{1:t-1}$ is stored. During training for the next task, while the feature extractor is updated, the classifier is trained using a classification loss on prototypes, perturbed with Gaussian noise, along with the current task data. This strategy enables the classifier to distinguish classes across different tasks, mitigating inter-task confusion and task-recency bias problems [20, 21, 22, 23, 55].
However, most of these prototype-rehearsal strategies rely on a strong feature distillation regularizer (see Eq. 5) to be effective, which inherently restricts plasticity. As a result, they are not suitable for Cold Start scenarios, where high plasticity is crucial to learning new tasks [29]. To address this limitation, we propose to decouple the feature-extractor learning phase, incorporating our EFM regularizer (Eq. 10), from the prototype rehearsal phase. Specifically, we introduce a prototype re-balancing training strategy, applied after training the feature extractor, which leverages both prototypes and current task data. This approach mitigates inter-task confusion and task-recency bias while preserving high plasticity, making it particularly well-suited for Cold Start scenarios.
IV-B Prototype Re-balancing Training
@@
where $\mathbf{X}\sim\mathcal{X}_{t}$ and $W_{t-1:t}$ represents the weights of the classifier associated with the latest classes. The cross-entropy loss is evaluated only on the current task classes, and is used to learn the features of the current task. $\mathcal{L}_{t}^{\text{EFM}}$ is the regularization loss based on Empirical Feature Matrix (see Eq. (10)) and is used to mitigate feature drift in the relevant directions for the previous task while allowing plasticity in the other directions.
@@
IV-C Prototype Drift Compensation via EFM
Fixed prototypes can become inaccurate over time due to drift from previous class representations. Yu et al. [24] showed that the drift in embedding networks can be estimated using current task data. They suggest estimating the drift of class means from previous-tasks by looking at feature drift in closely related classes from current task. They leveraged this drift to update class means in nearest-mean classification.
Let $p^{c}_{t-1}$ the prototype of the class $c$ at task $t-1$. Our aim is to update $p^{c}_{t-1}$ after the training of task $t$ to account for backbone drift. These updated prototypes are used for prototype rehearsal for training the model classifier. As proposed by SDC [24], we update the prototypes after the backbone is trained for each task as follows:
@@ IV-C Prototype Drift Compensation via EFM
where $\delta_{i}$ is the feature drift of samples $x_{i}$ between task $t-1$ and $t$.
We define the weights $w_{i}$ using the EFM (Figure 2b). In Section III-C we showed that the EFM induces a pseudo-metric in feature space, providing a second-order approximation of the KL-divergence due to feature drift. We extend the idea of feature drift estimation to softmax-based classifiers by weighting the overall drift of our class prototypes. Writing $p^{c}=p^{c}_{\text{t}-1}$ and $f_{t\text{-}1}(x_{i})=f(x_{i};\theta_{t-1})$ to simplify notation, our weights are defined as:
@@
where $E_{t\text{-}1}$ is the EFM after training task $t-1$. Eq. 15 assigns higher weights to the samples more closely aligned to prototypes in terms of probability distribution. Specifically, higher weights are assigned to samples whose softmax prediction matches that of the prototypes, indicating a strong similarity for the classifier.
IV-D Elastic Feature Consolidation++
To summarize, the training process begins with feature extractor training, where we optimize Eq. 12 to learn the new task while simultaneously regularize relevant directions from previous tasks (Figure 2a). Once the feature extractor is trained, we update the prototypes using the EFM (see Eqs. 1415 in Section IV-C), ensuring that representations from the previous task remain aligned with the current task drift (Figure 2b).
After updating the prototypes, we perform the prototype re-balancing phase to train the classifier (Eq. 13). Since the prototypes have already been aligned with the updated feature extractor, this guarantees that the classifier is trained with consistent representations (Figure 2c).
Finally, new prototypes and their corresponding class covariances are computed. The Empirical Feature Matrix $E_{t}$ is then calculated using the current task data and to be used as regularizer for the next incremental task (Figure 2d).
We call this procedure, detailed in Algorithm 1, Elastic Feature Consolidation++ (EFC++) to highlight its extension of our previous approach, Elastic Feature Consolidation (EFC), introduced in [29]. In the next section, we demonstrate why EFC++ is more effective in Cold Start scenarios and how it differs from and improves upon EFC.
@@ IV-D Elastic Feature Consolidation++
Algorithm 1 Elastic Feature Consolidation with Prototype Re-balancing
V EFC++ Mitigates the Prototype-Rehearsal Shortcomings of EFC
In this section, we first briefly review Elastic Feature Consolidation (EFC) [29]. We then highlight the limitations of the prototype-rehearsal strategy employed by EFC and, finally, explain why and how EFC++, introduced in this extension, mitigates these issues.
V-A Background: Elastic Feature Consolidation
Elastic Feature Consolidation (EFC) [29] is a combination of the empirical feature matrix loss $\mathcal{L}_{t}^{\text{EFM}}$ (Eq. 10), inhibiting drift in important directions for previous tasks, and an asymmetric cross-entropy loss (PR-ACE) balancing prototypes and new task samples during the current feature extractor training:
@@ V-A Background: Elastic Feature Consolidation
V-B Limitations of PR-ACE in Cold Start
@@
This misalignment is further compounded by the cross-entropy term $\mathcal{L}^{\text{ce}}_{t}(\tilde{\mathbf{P}}\cup f(\hat{\mathbf{X}},\theta_{t});W_{t})$ on the right-hand side of Equation 17, which induces an additional drift in the feature representation along the relevant directions for previous tasks identified by the EFM. This drift arises from the backward pass on $\hat{\mathbf{X}}$, which updates the weights $\theta_{t}$ and alters the feature representations. Because PR-ACE uniformly samples prototypes $\tilde{\mathbf{P}}$ and new data $\hat{\mathbf{X}}$ based on class labels, mini-batches contain a larger proportion of current task data in earlier tasks compared to the later tasks. Consequently, earlier tasks undergo stronger representation drift (see Figure 5 for drift analysis).
For illustration, consider the Cold Start scenario where each task introduces an equal number of classes. In the second task, PR-ACE forms batches containing 50% prototypes and 50% image data; in the third task, this shift to two-thirds prototypes and one-third current data. Because the cross-entropy loss is averaged over both prototypes and new data, the gradient updates on the feature extractor become less severe for later tasks. As more tasks are added, the drift in the feature representation along the relevant directions of the EFM diminishes. Consequently, prototypes from earlier tasks remain farther from the actual class means than those from more recent tasks, leading to forgetting. This effect is illustrated in Figure 4 (left), where in EFC the performance on the first five tasks declines faster than on the last five.
Managing the drift caused by PR-ACE is particularly crucial in the Cold Start scenario, where, during initial training, a strong backbone cannot be learned, making the representations more susceptible to change compared to the Warm Start scenario (as we will show in Section VIII-B). In the next section, we describe how EFC++ effectively manages this drift and provide empirical evidence of its efficacy.
V-C EFC++ addresses the PR-ACE limitations in Cold Start
@@
To validate this claim, we consider a model continuously trained using either EFC or EFC++ up to task $t-1$. We then measure the drift of the class means before and after training task $t$. Specifically, for each class $c\in\mathcal{C}_{t}$, we denote the feature representation of class $c$ prior to task $t$ by $\mu^{c}_{t-1}$ and after training task $t$ by $\mu^{c}_{t}$. These representations are obtained by averaging the feature representations of the data for class $c$. We define the feature drift of the class mean as the difference between these feature representations, specifically,
@@
$||\Delta\mu^{c}||_{E_{t}}=(\Delta\mu)^{\top}E_{t}\Delta\mu,$ (19)

### v3

@@ 3.1 Exemplar-free Class-Incremental Learning (EFCIL)
3.2 Weight Regularization and Feature Distillation
Kirckpatrick et al. ewc showed that using $\ell_{2}$ regularization to constrain weight drift reduces forgetting but does not leave enough plasticity for learning new tasks. Hence, they proposed Elastic Weight Consolidation (EWC), which relaxes the $\ell_{2}$ constraint with a quadratic constraint based on a diagonal approximation of the Empirical Fisher Information Matrix (E-FIM):
@@ 3.2 Weight Regularization and Feature Distillation
where $\theta_{t}^{*}$ is the model trained after task $t$. The E-FIM induces a pseudo-metric in parameter space amari encouraging parameters to stay in a low-error region for previous-task models:
@@
The main limitation of approaches of this type is need for approximations of the E-FIM (e.g. a diagonal assumption). This makes the computation of the E-FIM tractable, but in practice is unrealistic since it does not consider off-diagonal entries that represent influences of interactions between weights on the log-likelihood. To overcome these limitations, more recent EFCIL approaches pass ; ssre rely on feature distillation (FD), which scales better in the number of parameters:
[…]
@@
The isotropic regularizer using the $\ell_{2}$ distance is too harsh a constraint for learning new tasks podnet . We propose to use a pseudo-metric in feature space, induced by a matrix which we call the Empirical Feature Matrix (EFM), which constrains directions in feature space most important for previous tasks, while allowing more plasticity in other directions when learning new tasks. As shown in Figure 1, increasing plasticity is particularly relevant in Cold Start incremental learning since, in the initial task, a high-quality feature representation cannot be learned and methods must be plastic to adapt their backbone to new tasks.
3.3 The Empirical Feature Matrix
Figure 3: The regularizing effects of $E_{t}$ on the Cold Start CIFAR-100 - 10 and 20 step scenarios (see Section 6 for details on dataset settings). Left: Perturbing features in the principal directions of $E_{1}$ results in significant changes in classifier outputs (in blue), while perturbations in non-principal directions leave the outputs unchanged (in red). Middle: If we continue incremental learning up through task 3 and perturb features from all three tasks in the principal (solid lines) and non-principal (dashed lines) directions of $E_{3}$, we see that $E_{3}$ captures all important directions in feature space up through task 3. Right: At the end of training, we observe the same behavior: the last per-step accuracy (see Eq. 21), representing the average accuracy over all tasks after the last training session, decreases only when perturbed in directions of $E_{10}$ or $E_{20}$ relevant for previous tasks in the 10-step and 20-step scenarios, respectively.
@@ 3.3 The Empirical Feature Matrix
To gain additional insight into what this matrix measures, we highlight some (information) geometry aspects of $E_{t}$. The Empirical Fisher Information Matrix provides information about the information geometry of parameter space. In particular, it can be obtained as the second derivative of the KL-divergence for small weight perturbations naturalgradient ; datamatrix . Similarly, the Empirical Feature Matrix provides information about the information geometry of feature space. Since $E_{f_{t}(x)}$ is positive semi-definite, we can interpret it as a pseudo-metric in feature space and use it to understand which perturbations $\delta$ of features most affect the predicted probability:
@@
where $\mathbf{X}\sim\mathcal{X}_{t}$. In Eq. 10 we employ a damping term $\eta\in\mathbb{R}$ to constrain features to stay in a region where the quadratic second-order approximation of the KL-divergence of the log-likelihood in Eq. LABEL:kl_features remains valid Martens2012TrainingDA . However, to ensure that our regularizer does not degenerate into feature distillation, we must ensure that $\lambda_{\text{EFM}}\nu_{i}>\eta$, where $\nu_{i}$ are the non-zero eigenvalues of the EFM. Additional analysis of these constraints and the spectrum of the EFM are given in Section 8.3.
@@
Finally, the rightmost plot in Figure 3 illustrates the per-step accuracy at the end of the training, reflecting the overall continual learning performance (see Section 6 for a formal definition). As observed earlier, perturbing features along the principal directions of $E_{10}$ (or $E_{20}$) rapidly decreases performance, while perturbations in non-principal directions have no impact on the overall accuracy. This confirms that this behavior remains consistent across all tasks.
4 Elastic Feature Consolidation with Prototype Re-balancing
@@
4.1 Prototype Rehearsal for EFCIL
EFCIL suffers from the fact that the final classifier is never jointly trained on the all seen classes. At each task $t$, the classifier $W_{t-1}$ is extended to include $|\mathcal{C}_{t}|$ new outputs, resulting in a classifier with weights $W_{t}=[W_{t-1},W_{t-1:t}]$. Thus, if classifier $W_{t}$ is only trained on samples from the current classes $\mathcal{C}_{t}$, it becomes incapable of discriminating classes from previous tasks, leading to inter-task confusion. Additionally, the classifier is biased towards the newest classes, resulting in task-recency bias facil ; Wu_2019_CVPR .
A common exemplar-free solution is to replay only prototypes of previous classes generative_features ; pass ; ssre ; evanescent ; fetril ; il2a . At the end of task t1t-1, the prototypes $p^{c}_{t-1}$ are computed as the class means of deep features, and the set of prototypes $\mathcal{P}_{1:t-1}$ is stored. During training for the next task, while the feature extractor is updated, the classifier is trained using a classification loss on prototypes, perturbed with Gaussian noise, along with the current task data. This strategy enables the classifier to distinguish classes across different tasks, mitigating inter-task confusion and task-recency bias problems generative_features ; pass ; ssre ; evanescent ; il2a .
However, most of these prototype-rehearsal strategies rely on a strong feature distillation regularizer (see Eq. 5) to be effective, which inherently restricts plasticity. As a result, they are not suitable for Cold Start scenarios, where high plasticity is crucial to learning new tasks magistri2024elastic . To address this limitation, we propose to decouple the feature-extractor learning phase, incorporating our EFM regularizer (Eq. 10), from the prototype rehearsal phase. Specifically, we introduce a prototype re-balancing training strategy, applied after training the feature extractor, which leverages both prototypes and current task data. This approach mitigates inter-task confusion and task-recency bias while preserving high plasticity, making it particularly well-suited for Cold Start scenarios.
4.2 Prototype Re-balancing Training
@@
where $\mathbf{X}\sim\mathcal{X}_{t}$ and $W_{t-1:t}$ represents the weights of the classifier associated with the latest classes. The cross-entropy loss is evaluated only on the current task classes, and is used to learn the features of the current task. $\mathcal{L}_{t}^{\text{EFM}}$ is the regularization loss based on Empirical Feature Matrix (see Eq. 10) and is used to mitigate feature drift in the relevant directions for the previous task while allowing plasticity in the other directions.
@@
4.3 Prototype Drift Compensation via EFM
Fixed prototypes can become inaccurate over time due to drift from previous class representations. Yu et al. sdc showed that the drift in embedding networks can be estimated using current task data. They suggest estimating the drift of class means from previous-tasks by looking at feature drift in closely related classes from the current task. They leveraged this drift to update the class means in nearest-mean classification.
Let $p^{c}_{t-1}$ the prototype of the class $c$ at task t1t-1. Our aim is to update $p^{c}_{t-1}$ after the training of task $t$ to account for backbone drift. These updated prototypes are used for prototype rehearsal for training the model classifier. As proposed by SDC sdc , we update the prototypes after the backbone is trained for each task as follows:
@@ 4.3 Prototype Drift Compensation via EFM
where $\delta_{i}$ is the feature drift of samples $x_{i}$ between task t1t-1 and $t$.
We define the weights $w_{i}$ using the EFM (Figure 2b). In Section 3.3 we showed that the EFM induces a pseudo-metric in feature space, providing a second-order approximation of the KL-divergence due to feature drift. We extend the idea of feature drift estimation to softmax-based classifiers by weighting the overall drift of our class prototypes. Writing $p^{c}=p^{c}_{\text{t}-1}$ and $f_{t\text{-}1}(x_{i})=f(x_{i};\theta_{t-1})$ to simplify notation, our weights are defined as:
@@
where $E_{t\text{-}1}$ is the EFM after training task t1t-1. Eq. 15 assigns higher weights to the samples more closely aligned to prototypes in terms of probability distribution. Specifically, higher weights are assigned to samples whose softmax prediction matches that of the prototypes, indicating a strong similarity for the classifier.
4.4 Elastic Feature Consolidation++
In summary, the training process begins with feature extractor training, where we optimize Eq. 12 to learn the new task while simultaneously regularize relevant directions from previous tasks (Figure 2a). Once the feature extractor is trained, we update the prototypes using the EFM (see Eqs. 1415 in Section 4.3), ensuring that representations from the previous task remain aligned with the current task drift (Figure 2b).
After updating the prototypes, we perform the prototype re-balancing phase to train the classifier (Eq. 13). Since the prototypes have already been aligned with the updated feature extractor, this guarantees that the classifier is trained with consistent representations (Figure 2c). Finally, new prototypes and their corresponding class covariances are computed. The Empirical Feature Matrix $E_{t}$ is then calculated using the current task data and is used as regularizer for the next incremental task (Figure 2d).
We call this procedure, detailed in Algorithm 1, Elastic Feature Consolidation++ (EFC++) to highlight its extension of our previous approach, Elastic Feature Consolidation (EFC), introduced in magistri2024elastic . In the next section, we demonstrate why EFC++ is more effective in Cold Start scenarios and how it differs from and improves upon EFC.
@@ 4.4 Elastic Feature Consolidation++
Algorithm 1 EFC++: Elastic Feature Consolidation with Prototype Re-balancing
5 EFC++ Mitigates the Prototype-Rehearsal Shortcomings of EFC
In this section, we first briefly review Elastic Feature Consolidation (EFC) magistri2024elastic . We then highlight the limitations of the prototype-rehearsal strategy employed by EFC and, finally, explain why and how EFC++, introduced in this extension, mitigates these issues.
5.1 Background: Elastic Feature Consolidation
Elastic Feature Consolidation (EFC) magistri2024elastic is a combination of the empirical feature matrix loss $\mathcal{L}_{t}^{\text{EFM}}$ (Eq. 10), inhibiting drift in important directions for previous tasks, and an asymmetric cross-entropy loss (PR-ACE) balancing prototypes and new task samples during the current feature extractor training:
@@ 5.1 Background: Elastic Feature Consolidation
5.2 Limitations of PR-ACE in Cold Start
@@
This misalignment is further compounded by the cross-entropy term $\mathcal{L}^{\text{ce}}_{t}(\tilde{\mathbf{P}}\cup f(\hat{\mathbf{X}},\theta_{t});W_{t})$ on the right-hand side of Equation 17, which induces an additional drift in the feature representation along the relevant directions for previous tasks identified by the EFM. This drift arises from the backward pass on $\hat{\mathbf{X}}$, which updates the weights $\theta_{t}$ and alters the feature representations. Because PR-ACE uniformly samples prototypes $\tilde{\mathbf{P}}$ and new data $\hat{\mathbf{X}}$ based on class labels, mini-batches contain a larger proportion of current task data in earlier tasks compared to later tasks. Consequently, earlier tasks undergo stronger representation drift (see Figure 5 for drift analysis).
For illustration, consider the Cold Start scenario where each task introduces an equal number of classes. In the second task, PR-ACE forms batches containing 50% prototypes and 50% image data; in the third task, this shifts to two-thirds prototypes and one-third current data. Because the cross-entropy loss is averaged over both prototypes and new data, the gradient updates on the feature extractor become less severe for later tasks. As more tasks are added, the drift in the feature representation along the relevant directions of the EFM diminishes. Consequently, prototypes from earlier tasks remain farther from the actual class means than those from more recent tasks, leading to forgetting. This effect is illustrated in Figure 4 (left), where in EFC the performance on the first five tasks declines faster than on the last five.
Managing the drift caused by PR-ACE is particularly crucial in the Cold Start scenario, where, during initial training, a strong backbone cannot be learned, making the representations more susceptible to change compared to the Warm Start scenario (as we will show in Section 8.2). In the next section, we describe how EFC++ effectively manages this drift and provide empirical evidence of its efficacy.
5.3 EFC++ addresses the PR-ACE limitations in Cold Start
@@
To validate this claim, we consider a model continuously trained using either EFC or EFC++ up to task t1t-1. We then measure the drift of the class means before and after training task $t$. Specifically, for each class $c\in\mathcal{C}_{t}$, we denote the feature representation of class $c$ prior to task $t$ by $\mu^{c}_{t-1}$ and after training task $t$ by $\mu^{c}_{t}$. These representations are obtained by averaging the feature representations of the data for class $c$. We define the feature drift of the class mean as the difference between these feature representations, specifically,
@@
$\Delta\mu^{c}||_{E_{t}}=(\Delta\mu)^{\top}E_{t}\Delta\mu,$ (19)